OpenStudy (blackops2luvr):

Solve the equation for x. 3x = –4y

OpenStudy (shaik0124):

divide by 3 on both sides

OpenStudy (shaik0124):

u will get x

OpenStudy (blackops2luvr):

wait

OpenStudy (blackops2luvr):

A. x = -12 B. x = -12/y C. x = -3/4y D. x = -4y/3

OpenStudy (blackops2luvr):

ok......wat is it

OpenStudy (blackops2luvr):

sooo

OpenStudy (blackops2luvr):

ahhh....... u there?

OpenStudy (blackops2luvr):

@SolomonZelman

OpenStudy (blackops2luvr):

@Starr_DynastyT

OpenStudy (blackops2luvr):

@CGGURUMANJUNATH

OpenStudy (shaik0124):

d is answer

OpenStudy (blackops2luvr):

@hartnn

OpenStudy (blackops2luvr):

nope

OpenStudy (blackops2luvr):

@help.me

OpenStudy (solomonzelman):

Divide both sides by 3.

OpenStudy (shaik0124):

x=-4/3 y is the answer

OpenStudy (solomonzelman):

Why are we giving out plain answers, guys?

OpenStudy (blackops2luvr):

idk peeps do tht junk...not break up the prob......thats how people learn...not by giving answers

OpenStudy (cggurumanjunath):